OpenStudy (anonymous):

If a cube with side length 6 inches has its dimensions divided in half, what will be the volume of the new cube?

jimthompson5910 (jim_thompson5910):

Cut 6 in half to get ???

jimthompson5910 (jim_thompson5910):

now cube that

jimthompson5910 (jim_thompson5910):

3^3 = 3*3*3 = ??

OpenStudy (anonymous):

27

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

That's your answer since the volume is V = s^3 where V = volume s = side length

OpenStudy (anonymous):

yes its correct thank you..can u answer this? A fire truck has a ladder that can extend to 60 feet in length. The ladder can be safely raised to a maximum angle of 75° with the horizontal. Disregarding the height of the fire truck itself, which is closest to the maximum height that the ladder can safely reach?

jimthompson5910 (jim_thompson5910):

lets draw a pic first

jimthompson5910 (jim_thompson5910):

if we do, we might get something like this 

jimthompson5910 (jim_thompson5910):

our goal is to find h

jimthompson5910 (jim_thompson5910):

to tie h and 60 together (along with the angle 75 degrees), we use the sine function this is because sine = opposite/hypotenuse where opposite = h hypotenuse = 60

jimthompson5910 (jim_thompson5910):

so this means sin(75) = h/60

jimthompson5910 (jim_thompson5910):

solve for h, what do you get?

OpenStudy (anonymous):

second sine? or just sine?

jimthompson5910 (jim_thompson5910):

what do you mean by second sine

OpenStudy (anonymous):

on the scientific calculator if u r finding the angle you have to press second then sin to find the correct answer

jimthompson5910 (jim_thompson5910):

ah i see what you meant

OpenStudy (anonymous):

but now i see that the angel is given

jimthompson5910 (jim_thompson5910):

that's known as inverse sine or arcsine

OpenStudy (anonymous):

angle*

jimthompson5910 (jim_thompson5910):

if you have sin(x) = 1/2 for instance and you wanted to solve for x, you would apply the arcsine to both sides to get x = arcsin(1/2)

jimthompson5910 (jim_thompson5910):

to evaluate arcsin(1/2), you would type "2nd", "sin", "1", "/", "2", "enter"

OpenStudy (anonymous):

makes sense

jimthompson5910 (jim_thompson5910):

in our case, we don't need to use arcsine because we dont need to solve for the angle so we just isolate h like this sin(75) = h/60 60*sin(75) = h h = 60*sin(75) h = 60*0.9659258 ... use a calculator here (make sure you're in degree mode) h = 57.955548

OpenStudy (anonymous):

yay its correct

jimthompson5910 (jim_thompson5910):

lol that's good

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome